OpenStudy (anonymous):

Solve the following equation. One point is given for showing steps and one point given for the correct solution. x/2 + x/3 = 5

OpenStudy (campbell_st):

multiply every term by 6 \[\frac{x}{2} \times6 +\frac{x}{3}\times6 = 5\times6\] then simplifying \[3x + 2x = 30\] I'll let you finish the solution

OpenStudy (anonymous):

5x=30?

OpenStudy (campbell_st):

correct

OpenStudy (anonymous):

Yes! thank you for the help, but where did you get the 6 from?

OpenStudy (campbell_st):

to eliminate the fractions I multplied by 6, the common denominator.... 2 x 3 or the question can be solved as x/2+x/3 = 5 5x/6 = 5

OpenStudy (anonymous):

Oh ok ,thank you.